OpenStudy (anonymous):

b. Find the angle A for the triangle ABC given that: AB = 14mm BC = 29mm B = 90° Has to be done using trigonometric functions

OpenStudy (anonymous):



OpenStudy (anonymous):

What formula do I use?

OpenStudy (anonymous):

Taking BC as the perpendicular and AB as the base; \[\tan \theta =\frac{ perpedicular }{ base }\] \[\tan \theta= \frac{ 29 }{ 14 }\]

OpenStudy (anonymous):

I got 64.23 degrees (2dp) as my final answer.

OpenStudy (anonymous):

Yep. Correct.